OpenStudy (bananas1234):

Solve using perfect square factoring patterns. x^2 – 12x + 36 = 4

OpenStudy (anonymous):

(x-6)^2 -6 multiply this out =x^2-12x+36-6 =original but how did we get that you look at the 12x term take half put in the bracket (x-6)^2 multiply it out see what you have added and subtract it to balance if you add you must subtract to balance it takes longer to explain than do

OpenStudy (anonymous):

x^2 – 12x + 36 = 4 implies (x-6)^2=2^2 implies x-6=2 or -2 implies x=8 or x=4

OpenStudy (anonymous):

@Bananas1234

OpenStudy (anonymous):

x^2 – 12x + 36 = 4 implies (x-6)^2=2^2 implies (x-6)^2-2^2=0 implies (x-6+2)(x-6-2)=0 implies (x-4)(x-8)=0 implies x=4,8

OpenStudy (anonymous):

@Bananas1234